# Distribution-Free Calibration of Statistical Confidence Sets

### v1

@@ Appendix B Algorithm for tuning $M$
Data: Validation grid size $B_{\text{tune}}$; grid of $M$’s between 00 to $K$ $M_{\text{grid}}$; fitted TRUST++; number of simulated statistics $n_{\text{sim}}$ in each grid point
@@ C.2 Section 3 - Theoretical Results
$\theta_{1},\tau(\mathbf{X}_{1},\theta_{1})),\ldots,(\theta_{B},\tau(\mathbf{X}_{B},\theta_{B})),$
@@
Proof [Proof of Theorem 2] This is a straightforward application of conformal prediction (see Angelopoulos et al. (2023)). The complete proof can be found in detail in (Cabezas et al., 2024, Theorem 2).
@@
[…]
@@
[…]
@@
This implies that, as long as our approximation $\widehat{H}_{B}$ closely matches $H$, our partition-based estimate will serve as a reliable approximation of the test statistics distribution. In particular, as discussed in Section 2.1.1, let $\widehat{C}_{\theta,B}=\widehat{H}^{-1}_{B}(\alpha|\theta)$ denote the $\alpha$-quantile of the values $\{\tau(\mathbf{X}_{b},\theta_{b}):b\in I_{A(\theta)}\}$. Then,
@@
Since $H(t|\theta)$ is constant (not random), it follows that
@@
Following (Meinshausen and Ridgeway, 2006, Assumption 3), we assume there exists $0.50<\gamma<$ 0.5 such that $\mathbb{P}(\theta^{\prime}\in A(\theta))\geq\gamma B$. This assumption is reasonable, as it implies that the partitions (i.e., the leaves) generated by the tree-based estimators are uniformly well-populated. In other words, each leaf contains a sufficient number of observations, ensuring that the coverage properties hold consistently across the partitions. Therefore,

### v2

@@ Appendix B Algorithm for tuning $M$
Data: Validation grid size $B_{\text{tune}}$; grid of $M$’s between 0 to $K$ $M_{\text{grid}}$; fitted TRUST++; number of simulated statistics $n_{\text{sim}}$ in each grid point
@@ C.2 Section 3 - Theoretical Results
$(\theta_{1},\tau(\mathbf{X}_{1},\theta_{1})),\ldots,(\theta_{B},\tau(\mathbf{X}_{B},\theta_{B})),$
@@
Proof [Proof of Theorem 2] This is a straightforward application of conformal prediction (see Angelopoulos et al. (2023)). The complete proof can be found in detail in (Cabezas et al., 2025, Theorem 2).
@@
Let $H(t|\theta)=\mathbb{P}(\tau(\mathbf{X},\theta)\leq t|\theta)$ and $A$ be a measurable set. Suppose $p(x,\theta)$ is the joint PDF of $(\mathbf{X},\theta)$, satisfying $\int p(x,\theta)\,dx=r(\theta)>0$. Then we have
@@
$=c^{-1}\int H(t|\theta)\mathbb{I}[\theta\in A]r(\theta)d\theta$
$=c^{-1}\int\mathbb{P}(\tau(\mathbf{X},\theta)\leq t|\theta)\mathbb{I}[\theta\in A]r(\theta)d\theta$
$=c^{-1}\int\left(\frac{1}{r(\theta)}\int\mathbb{I}[\tau(x,\theta)\leq t]p(x,\theta)dx\right)\mathbb{I}[\theta\in A]r(\theta)d\theta$
$=c^{-1}\int\int\mathbb{I}[\tau(x,\theta)\leq t]\mathbb{I}[\theta\in A]p(x,\theta)dxd\theta$
@@
This implies that, as long as our approximation $\widehat{H}_{B}$ closely matches $H$, our partition-based estimate will serve as a reliable approximation of the test statistics distribution. In particular, as discussed in Section 2.1.1, let $\widehat{C}_{\theta,B}=\widehat{H}^{-1}_{B}(\alpha|\theta)$ denote the adjusted $\alpha$-quantile of the values $\{\tau(\mathbf{X}_{b},\theta_{b}):b\in I_{A(\theta)}\}$. Then,
@@
Since $H(t|\theta)$ is constant, it follows that
@@
Following (Meinshausen and Ridgeway, 2006, Assumption 3), we assume there exists $0<\gamma<0.5$ such that $\mathbb{P}(\theta^{\prime}\in A(\theta))\geq\gamma B$. This assumption is reasonable, as it implies that the partitions (i.e., the leaves) generated by the tree-based estimators are uniformly well-populated. In other words, each leaf contains a sufficient number of observations, ensuring that the coverage properties hold consistently across the partitions. Therefore,